Write the slope-intercept equation for the line that passes through (4, -9) and has slope -3/4. Please show all of your work
Answer by mananth(16946)

m= -0.75 - 3/ 4 12/29/99

Plug value of the slope and point ( 4,-9 ) in
Y = m x + b
-9.00 = -3 + b
b= -9-(-3)
b= -6
So the equation will be
Y = -3/4 x -6
m.ananth@hotmail.ca